OpenStudy (amonoconnor):

What is the derivative of: x^2/8 ?

OpenStudy (luigi0210):

\[\LARGE \frac{(8)(x^2)'-(8)'(x^2)}{(8)^2}\] If you know the quotient rule, you should know this

OpenStudy (amonoconnor):

16x/64 ?

OpenStudy (amonoconnor):

No wait...

OpenStudy (luigi0210):

No, you're right, now simplify

OpenStudy (amonoconnor):

(16x - x^2)/ 64 ?

OpenStudy (amonoconnor):

Oh...

OpenStudy (luigi0210):

Remember, (8)'=0 and 0*x^2=0 So 16x-0= 16x/64 and 64 goes into 16 4 times, so x/4 if I'm not mistaken :)

OpenStudy (amonoconnor):

Ahh yes! No variable = 0, when derived. Okay. I get it.... The neurons connected! :) Thank you thank you thank you!!

OpenStudy (luigi0210):

No problem :)